A manager of a store that sells and installs spas wants to prepare a forecast for January, February, and March of next year. Her forecasts are a combination of trend and seasonality. She uses the following equation to estimate the trend component of monthly demand: Ft = 60 + 4t, where t = 0 in June of last year. Seasonal relatives are .89 for January, .95 for February, and 1.11 for March. What demands should she predict? (Round your answers to 2 decimal places.)
Answer #1
last year jan
feb
mar
apr
may
last year jun 0
jul 1
aug 2
sep 3
oct 4
nov 5
dec 6
this year jan 7
feb 8
mar 9
apr 10
may 11
jun 12
jul 13
aug 14
sep 15
oct 16
nov 17
dec 18
We need this month jan next year jan 19
We need this feb next year feb 20
We need this month of march next year mar 21

---- ---  

--

--

This is best done in 2 stages:

Stage 1:

Compute the Trend = Yt

We need to find Y Jan Next Year, Y Feb Next Year, and Y March Next year which is just 19,20, and 21 ( see the excel table – it is just a simple sequential counting of months from June last year at 0 to Jan next year = one and a half years, hence 19 months.)

Formula for trend is already given as 60 + 4t

The value of t for the 3 months is found by simply adding as follows:

Ft = 60 + 4 * t

Ft for Jan = Y Jan Next Year = 60 + 4 * 19 = 136

Ft for Feb = Y Feb Next Year = 60 + 4 * 20 = 140

Ft for Mar = Y Mar Next Year = 60 + 4 * 21 = 144

Stage 2:

Compute the forecast = trend * seasonal relative

The seasonal relative for each month is already given in the question as 0.89, 0.95, and 1.11 for the 3 months respectively.

We found the 3 trends above as 136, 140, and 144 for the 3 months respectively.

Hence January fore cast = 0.89 * 136 = 121.04

February fore cast = 0.95 * 140 = 133

March fore cast = 1.11 * 144 = 159.79994

--

--